OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10)and (-4, 0)? -2/5 5/2 -2/5 2/5

OpenStudy (anonymous):

change in x/change in y

OpenStudy (inkyvoyd):

@Snipa420 , that's actually wrong.

OpenStudy (anonymous):

change in y/change in x. my bad

OpenStudy (inkyvoyd):

slope=rise/run=(change in y)/(change in x)

OpenStudy (anonymous):

So m=(y2-y1)/(x2-x1)

OpenStudy (inkyvoyd):

What @Snipa420 , with the points (x1,y1) and (x2,y2)

OpenStudy (mertsj):

The correct answer is not in your list.

OpenStudy (anonymous):

It is -5/2

OpenStudy (inkyvoyd):

(-10-0)/(0--4)=(-10)/(4)=-5/2